OpenStudy (anonymous):

From the deck of cards shown, two cards are drawn with replacement. The probability of drawing a yellow followed by a yellow is 1 out of ____.

OpenStudy (anonymous):



OpenStudy (raden):

P(Y and Y) = P(Y) * P(Y|Y) = 1/10 * 1/10 = 1/... ?